Suppose the market demand curve for a good is Qd=1000-10p and the market supply curve is given by Qs=100+20p required; calculate the equilibrium price and the equilibrium quantity

1
Expert's answer
2021-11-09T17:51:59-0500

At equilibrium price, both "Qd = Qs",

Therefore, "Qd=1000-10p=Qs=100+20p \\
1000-10p =100 + 20p\\
1000-100 = 20p+10p\\
900=30p\\
p=30"

Equilibrium price = 30

"Qd=1000-10p\\
Qd = 1000-10(30)\\
Qd = 700"

Equilibrium quantity = 700